OpenStudy (anonymous):

i need some help with this one Solve. x2 + 5x – 6 = 0 A. {–2, –3} B. {–1, 6} C. {1, –6} D. {2, 3}

OpenStudy (anonymous):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

@thebeartellsall

Nnesha (nnesha):

leading coefficient is one so i would apply *headphone method* find two number if you multiply them you should get product of AC and if you add or subtract them you should get middle term b \[\huge\rm Ax^2 +Bx+C=0\]

OpenStudy (anonymous):

ok i'll try that

OpenStudy (anonymous):

I"m getting (-2,-3)

Nnesha (nnesha):

-2 times -3 = 6 -2 -3 =-5 when you multiply them you should get product of AC which is what in this equation ?

OpenStudy (anonymous):

-2-3

Nnesha (nnesha):

nope

Nnesha (nnesha):

what is a and c in that equation ?

OpenStudy (anonymous):

C

Nnesha (nnesha):

\[\bf Ax^2 +Bx+C\] leading coefficient is a and constant term is c

OpenStudy (anonymous):

answer is C @speartonion

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

@drkid15 you are not allowed to give direct answer